1. On a map, the distance between the two towns is 2.6 cm. The scale of the map is 1 cm: 50 km. What is the actual distance between the two towns ? and whats the process to get that answer
1
Expert's answer
2022-01-05T17:55:29-0500

To get the answer we should solve the proportion: "\frac{50\;km}{1\;cm}=\frac{x\;km}{2.6\;cm}."

Thus, "x=2.6*50=130 \;km."